OpenStudy (anonymous):

Katalin drove 300 miles on her vacation. She drove an average of 1.9 times faster on the second 150 miles of her trip than she did on the first 150 miles of her trip. Which expression represents the time she spent driving? Let x = her speed on the first half of the trip.

OpenStudy (mathstudent55):

t1 is the time the first 150 miles took. t2 is the time the second 150 miles took. average speed = distance/time 150 mi/t1 = x first 150 mi 150 mi/t2 = 1.9x second 150 mi t1 = 150 mi/x t2 = 150 mi/1.9x Total time, t = t1 + t2 = 150 mi/x + 150 mi/(1.9x)

OpenStudy (mathstudent55):

Do you have choices?

OpenStudy (anonymous):

457.89/x 228.95/x 228.95x 435/x

OpenStudy (mathstudent55):

The answer is 150/x + 150/1.9x = 150/x * 1.9/1.9 + 150/1.9x = 285/1.9x + 150/1.9x = 435/1.9x = 228.95/x

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

wlcm